### v3

@@ III-B2 Limitations
Environment
Crowded places
[59, 65, 68, 95]
Dynamic object
[59, 90, 68, 73, 109, 111]
Sensor
Fast motion
[72]
Illumination
[61, 71, 72, 74, 81]
III-C1 Environmental complexities
@@
Crowded scenarios Navigating crowded environments presents significant challenges for the visually impaired, leading to increased collision risks and difficulties in maintaining personal space. The absence of visual information makes it difficult to measure the distance, perceive crowd density, and locate landmarks or places of interest.
@@
Navigation in crowded environments also poses challenges for assistive technology. For example, in assistive systems that operate based on SLAM, the presence of numerous dynamic elements, such as moving individuals and objects, introduces ambiguity into feature detection and tracking, leading to difficulties in accurately estimating the pose of the user and structure of the environment. The dynamic nature of crowds also hinders loop closure detection, disrupts map consistency, and contributes to drifts. Moreover, the lack of distinct visual landmarks in crowded scenes represents a reliable localization challenge that potentially reduces the robustness and accuracy of the overall SLAM system. Addressing these challenges requires the development of new approaches to address the complexity of such environments effectively. Several studies have examined this issue.
@@
[59] presented a guide mobile robot engineered for the complexities of navigating different environments while considering dynamic objects and human presence. The robot could handle crowded environments with multiple dynamic objects. To accomplish this, the robot leveraged a spatial risk map, a tool that evaluates potential object-occupied spaces, to chart a path that effectively minimizes disruptions. This paper presents experiments in which a robot successfully guided a user through the passage of multiple objects and people.
@@
In another study, [65] introduced an egocentric human trajectory forecasting model designed for navigation in crowded environments. By leveraging a wearable camera, the model captures detailed surroundings and the user’s past trajectory to predict their future. The trajectories obtained using ORB-SLAM3 were used to train the egocentric human trajectory forecasting model, which predicted the wearer’s future trajectory based on the contextual cues captured by the wearable camera and the wearer’s past trajectory.
@@
Additionally, [68] addressed challenges in crowded environments using a combination of SLAM and Ultra-Wideband (UWB) positioning. However, the SLAM algorithm was found to be less effective in environments with dynamic obstacles such as pedestrians. The algorithm finds features of dynamic obstacles moving along with the robot as the assistive device, and thus it was misled that the robot did not move at all. However, using UWB positioning mitigates this issue.
@@
In [95], the method addressed crowded environments by recognizing and predicting people’s behavior while anticipating the collision risk. The system advises users to adjust their walking speed (on-path mode) or to choose alternative routes (off-path mode). This involves comparing the 3D point cloud map to the real-time LiDAR and IMU sensor data. The system then predicts the future position and velocity of the user in order to avoid collisions.
@@
Dynamic objects Several publications have not directly addressed the challenge of crowded environments, and have only focused on dealing with the presence of dynamic objects within the scene. In the system proposed by [90], dynamic objects can be identified, and average depth information can be provided to the user. When a dynamic object belongs to a predefined class such as a person, it can also be tracked between frames in the SLAM pipeline. The system is capable of identifying and tracking dynamic objects after ego-motion estimation to obtain average depth information. Subsequently, it can estimate the poses and speeds of these tracked dynamic objects and relay this information to the users through acoustic feedback. Depth information helps users maintain social distancing in public indoor environments such as shopping malls. To address the challenge of dynamic objects, [73] proposed a new method called visual simultaneous localization and mapping for moving person tracking (VSLAMMPT). This method was designed to handle dynamic environments in which objects constantly move. The system also uses the expected error reduction with active-semi-supervised learning (EER–ASSL)-based person detection to eliminate noisy samples in dynamic environments. This aids in accurate detection and avoidance of dynamic obstacles.
III-C2 Sensor-related
@@
Changes in lighting condition Lighting changes pose a hurdle to visual SLAM systems. Illumination variations alter the visual features, interfere with accurate detection and matching across frames, impact pose estimation, and map-building robustness. SLAM relies on distinctive features for operation; however, lighting changes introduce ambiguities, noise, and errors, which affect accuracy. Overcoming this challenge requires robust algorithms for dynamic lighting to ensure stable and precise localization and mapping.
@@
The method proposed in [71] tackles the challenge of illumination changes using a deep descriptor network called a Dual Desc, which is designed to be robust against various appearance variations including illuminance changes. The network uses multimodal images (RGB, Infrared, and Depth) to generate robust attentive global descriptors and local features. These descriptors were used to retrieve coarse candidates of query images, and 2D local features, along with a 3D sparse point cloud, were used for geometric verification to select the optimal results from the retrieved candidates. The authors mentioned that their dataset included images captured at different times of the day, resulting in illumination changes between the query and database images. Despite these changes, the proposed method achieved satisfactory localization results.
@@
The authors of [72] evaluated the influence of lighting conditions on the performance of their novel localization method. The authors captured training images during the day and test images at night, and simulated changes in lighting conditions by switching some of the lights off in locations without windows. The results showed that changes in the lighting conditions had a minor impact on the proposed method. [61] mentioned that the proposed localization scheme was verified in a typical office building environment with dramatically changing lighting conditions throughout the day; however, it does not provide detailed results or discussion on how changing lighting conditions affect the performance of the system.
@@
The method proposed in [81] addressed changes in illumination as a challenge using the COLD and IDOL datasets, which were recorded under different weather and illumination conditions (cloudy, night, and sunny) using different mobile platforms and camera setups. These datasets were used to evaluate the strength of the localization and recognition algorithms with respect to the variations caused by human activities and changes in illumination conditions. The study also mentioned the use of Histogram of Oriented Gradients (HOG) for feature extraction, which provides preferable invariant results for lighting or shadowing.
@@
Motion blur Blurred images in visual SLAM can lead to inaccuracies in feature detection and matching, causing issues with pose estimation, map-building, loop closure, and visual odometry. These inaccuracies can also impact depth measurements and map quality. To address this issue, strategies such as using high-frame-rate sensors, incorporating IMUs, and employing motion deblurring techniques can be employed to improve the accuracy of localization and mapping in SLAM systems. Motion blur can be caused by fast or sudden movements of the user during navigation, which can affect localization performance. [72] studied this challenge and evaluated the robustness of the localization methods. They captured 2316 blurred images on the testing day. The results showed that the proposed method performed poorly in this experiment, indicating that fast motion or sudden changes in user movement can pose a challenge to the system. The reason for this poor performance was that the object detection scores did not exceed the threshold during the experiment.
@@
This section discusses how the approaches proposed by the studies included in our SLR have the potential to improve navigation for BVI people. These studies focused on diverse attributes, such as accurate pose estimation, semantic mapping, sensor fusion, and algorithmic innovations, to improve the quality of BVI navigation. Table XIV presents the categorization of attributes that contribute to enhancing the mobility and navigation of visually impaired individuals.
@@
TABLE XIV: Attributes of SLAM-based navigation systems that contribute to enhancing BVI navigation, along with referenced papers emphasizing each feature.
@@
Using SLAM algorithms, these solutions accurately estimate the position and orientation of visually impaired users. Precise localization is essential for visually impaired navigation systems in order to ensure accurate real-time guidance, obstacle avoidance, and spatial awareness, ultimately enhancing independent and safe mobility.
[109, 97, 95, 93, 90, 82, 78, 77, 76, 75, 73, 70, 69, 64, 89, 67]
Real-time assistance
@@
This feature ensures that BVI users receive immediate feedback about their environment. Therefore, they can be guided to navigate safely and efficiently, thereby enhancing their overall mobility and independence.
[85, 57, 92, 82, 77, 69, 66, 67, 88]
Semantic mapping
@@
Semantic mapping involves creating detailed environmental representations that go beyond geometric data and enables visually impaired users to navigate with a deeper understanding of their surroundings.
[100, 97, 66, 72, 84]
Both indoor and outdoor navigation
@@
Systems that serve both indoor and outdoor settings allow users to transition seamlessly between different environments while receiving consistent support.
[85, 108, 67, 74, 111]
Innovative algorithms
@@
Innovative algorithms lead to advancements in navigation techniques. These approaches contribute to a more efficient and effective navigation, ultimately improving the overall experience of the visually impaired.
[77, 66, 106, 78, 91, 73]
Robotic navigation
@@
These solutions employ robotic systems, such as robots, smart canes, and suitcases, which are equipped with sensors to assist BVI in navigating their environment.
[112, 107, 78, 70, 64, 89, 67, 59, 68, 87]
Smartphone based

### v4

@@ III-C1 Technical and methodological challenges
Environment Crowded places [104, 59, 65, 68, 95]
Dynamic object [100, 71, 59, 90, 68, 73, 109, 111]
Sensor Fast motion [72]
Illumination [61, 71, 72, 74, 81]
Environmental complexities
@@
Crowded scenarios Navigating crowded environments presents significant challenges for the visually impaired, leading to increased collision risks and difficulties in maintaining personal spaces. The absence of visual information makes it difficult to measure distance, perceive crowd density, and locate landmarks or places of interest.
@@
Navigation in crowded environments also poses challenges for assistive technologies. For example, in assistive systems that operate based on SLAM, the presence of numerous dynamic elements, such as moving individuals and objects, introduces ambiguity into feature detection and tracking, leading to difficulties in accurately estimating the pose of the user and structure of the environment. The dynamic nature of crowds also hinders loop closure detection, disrupts map consistency, and contributes to drift. Moreover, the lack of distinct visual landmarks in crowded scenes represents a reliable localization challenge, which potentially reduces the robustness and accuracy of the overall SLAM system. Addressing these challenges requires the development of new approaches to address the complexity of such environments effectively. Several studies have investigated this issue.
@@
[59] presented a guide mobile robot engineered for the complexities of navigating different environments while considering dynamic objects and human presence. The robot could handle crowded environments with multiple dynamic objects. To accomplish this, the robot leveraged a spatial risk map, which is a tool that evaluates potential object-occupied spaces, to chart a path that effectively minimizes disruptions. This study presents experiments in which a robot successfully guided a user through the passage of multiple objects and people.
@@
In another study, [65] introduced an egocentric human trajectory forecasting model that was designed for navigation in crowded environments. By leveraging a wearable camera, the model captures detailed surroundings and the user’s past trajectory to predict their future. The trajectories obtained using ORB-SLAM3 were used to train the egocentric human trajectory forecasting model, which predicted the wearer’s future trajectory based on contextual cues captured by the wearable camera and the wearer’s past trajectory.
@@
In addition, [68] addressed challenges in crowded environments using a combination of SLAM and Ultra-Wideband (UWB) positioning. However, the SLAM algorithm was found to be less effective in environments with dynamic obstacles such as pedestrians. The algorithm finds features of dynamic obstacles moving along with the robot as the assistive device, and thus, it was misled that the robot did not move at all. However, using UWB positioning mitigates this issue.
@@
In [95], the method addressed crowded environments by recognizing and predicting people’s behavior while anticipating the collision risk. The system advises users to adjust their walking speed (on-path mode) or to choose alternative routes (off-path mode). This involves comparing the 3D point cloud map to real-time LiDAR and IMU sensor data. The system then predicts the future position and velocity of the user in order to avoid collisions.
An intelligent autonomous scooter was developed in [104] for navigating environments with small safety margins and highly dynamic pedestrian traffic such as sidewalks with numerous obstacles and pedestrians. The authors proposed a hybrid mapping solution that combines far-field and near-field mapping to navigate through dynamic environments. This approach utilizes sensor fusion to adapt dynamically to complex and cluttered environments.
@@
Dynamic objects Several publications have not directly addressed the challenge of crowded environments, and have only focused on dealing with the presence of dynamic objects within the scene. In the system proposed by [90], dynamic objects can be identified, and average depth information can be provided to the user. When a dynamic object belongs to a predefined class such as a person, it can also be tracked between frames in the SLAM pipeline. The system is capable of identifying and tracking dynamic objects after ego-motion estimation to obtain average depth information. Subsequently, it can estimate the poses and speeds of these tracked dynamic objects and relay this information to the users through acoustic feedback. Depth information helps users maintain social distancing in public indoor environments such as shopping malls.
To address the challenge of dynamic objects, [73] proposed a new method called visual simultaneous localization and mapping for moving person tracking (VSLAMMPT). This method was designed to handle dynamic environments in which objects constantly move. The system also uses expected error reduction with active-semi-supervised learning (EER–ASSL)-based person detection to eliminate noisy samples in dynamic environments. This aids in accurate detection and avoidance of dynamic obstacles.
[100] utilized YOLOv3 to detect common objects in a corridor, including people, which were identified as obstacles. The system sends information about obstacles to users every five seconds when the distance between the user and obstacle is less than 10 m. For example, it may notify the user, ”A person is located 2.8 meters ahead.”
Sensor-related challenges
@@
Changes in lighting condition Lighting changes pose a hurdle to visual SLAM systems. Illumination variations alter the visual features, interfere with accurate detection and matching across frames, impact pose estimation, and map-building robustness. SLAM relies on distinctive features for operation; however, lighting changes introduce ambiguities, noise, and errors, which affect the accuracy. Overcoming this challenge requires robust algorithms for dynamic lighting to ensure stable and precise localization and mapping.
@@
The method proposed in [71] tackles the challenge of illumination changes using a deep descriptor network called a Dual Desc, which is designed to be robust against various appearance variations including illuminance changes. The network used multimodal images (RGB, Infrared, and Depth) to generate robust attentive global descriptors and local features. These descriptors were used to retrieve coarse candidates from query images, and 2D local features, along with a 3D sparse point cloud, were used for geometric verification to select the optimal results from the retrieved candidates. The authors mentioned that their dataset included images captured at different times of the day, which resulted in illumination changes between the query and database images. Despite these changes, the proposed method achieved satisfactory localization results.
@@
The authors of [72] evaluated the influence of lighting conditions on the performance of their novel localization method. The authors captured training images during the day and test images at night and simulated changes in lighting conditions by switching some of the lights off in locations without windows. The results showed that changes in the lighting conditions had a minor impact on the proposed method.
[61] mentioned that the proposed localization scheme was verified in a typical office building environment with dramatically changing lighting conditions throughout the day; however, it does not provide detailed results or discussion on how changing lighting conditions affect the performance of the system.
@@
The method proposed by [81] addressed changes in illumination as a challenge using the COLD and IDOL datasets, which were recorded under different weather and illumination conditions (cloudy, night, and sunny) using different mobile platforms and camera setups. These datasets were used to evaluate the strength of the localization and recognition algorithms with respect to the variations caused by human activities and changes in illumination conditions. The study also mentioned the use of Histogram of Oriented Gradients (HOG) for feature extraction, which provides preferable invariant results for lighting and shadowing.
@@
Motion blur Blurred images in visual SLAM can lead to inaccuracies in feature detection and matching, causing issues with pose estimation, map building, loop closure, and visual odometry. These inaccuracies can also impact depth measurements and map quality. To address this issue, strategies such as using high-frame-rate sensors, incorporating IMUs, and employing motion-deblurring techniques can be employed to improve the accuracy of localization and mapping in SLAM systems.
Motion blur can be caused by fast or sudden movements of the user during navigation, which can affect localization performance. [72] studied this challenge and evaluated the robustness of the localization methods. They captured 2316 blurred images on the testing day. The results show that the proposed method performed poorly in this experiment, indicating that fast motion or sudden changes in user movement can pose a challenge to the system. The reason for this poor performance is that the object detection scores did not exceed the threshold during the experiment.
III-C2 Practical challenges and considerations
In addition to the technical and methodological aspects, we recognized the importance of practical challenges considerations that can significantly affect the usability and adoption of SLAM-based assistive systems. Therefore, we included an evaluation of the practical challenges and operational efficiency, as summarized in Tables XXIV and XXV. The information in these tables has either been directly extracted from the article or can be easily inferred from the article’s text. These tables provide information on user-friendliness, cost-efficiency, weight, comfort for extended use, adjustable fit, fatigue mitigation, and portability of the assistive tools described in the reviewed studies. For instance, while smartphones and lightweight devices such as eyeglasses-mounted sensors [74] and ARCore-supported smartphones with haptic gloves [75] are generally well received because of their high portability and ease of use, heavier devices such as guiding robots [68] and rolling suitcase-shaped device [95] are noted to cause user fatigue over extended periods. The augmented cane [67], although found to improve confidence and workload for novice and expert users, also faced usability challenges owing to its weight. Clear instructions and easy learning curves, as seen in electronic glasses with haptic modules [57], play a significant role in enhancing user satisfaction. However, the cost efficiency of these technologies varies, with some solutions being more affordable and accessible.
TABLE XXIV: Practical challenges and operational efficiency - Part I.
Ref. Assistive tool User-friendliness Cost-efficient Weight Comfort for extended use Adjustable fit Fatigue mitigation Portability
Clear instructions Easy to learn
[56] Smartphone $\checkmark$ $\checkmark$
[57] Electronic glasses and leg-mounted haptic modules $\checkmark$ $\checkmark$ Light $\checkmark$ $\checkmark$ $\checkmark$ High
[58] Microsoft Hololens2 $\checkmark$ Light $\checkmark$ High
[59] Wheeled guide mobile robot $\checkmark$ Light $\checkmark$ Easy-to-hold handle Moderate
[60] Android application $\checkmark$ $\checkmark$ Light $\checkmark$ $\checkmark$ $\checkmark$ High
[61] Not a user-based evaluation: only a technical test.
[62] Helmet; white cane $\checkmark$ Light $\checkmark$ Moderate
[63] Smartphone $\checkmark$ Need time to learn interpreting tactile signals
[64] Smart cane $\checkmark$ Light High
[65] No prototype implemented; the proposed approach for trajectory forecasting was tested with a robot.
[66] A forehead-mounted camera, an earphone, a computing resource bag Not a user-based evaluation: only a technical test performed. $\checkmark$ Heavy Heavyweight Heavyweight Moderate
[67] Augmented cane $\checkmark$ $\checkmark$ $\checkmark$ Heavy (1kg) Heavyweight $\checkmark$ Heavyweight Moderate
[68] Guiding robot $\checkmark$ $\checkmark$ USD 17000 Heavy (25kg) Comfortable handle feedback; some users noted slewing and speed change discomfort. The robot is relatively large at 41x43x25 $\mathrm{cm}^{3}$.
[69] Head-mounted camera Not a user-based evaluation: only a technical test performed. $\checkmark$ Light
[70] Smart cane Not a user-based evaluation: only a technical test performed.
[71] Auxiliary glasses Not a user-based evaluation: only a technical test performed.
[72] No prototype implemented: Only a technical test performed.
[73] Smart eyeglasses Limited instructions $\checkmark$ High
[74] Eyeglasses-mounted sensors + smartphone $\checkmark$ Participants trained in 10 minutes $\checkmark$ Light Prolonged beeping may cause discomfort. Glasses weight strains the nose. High
[75] ARCore-supported smartphone + haptic gloves $\checkmark$ A 5-minute tutorial $\checkmark$ High
[76] Optical see-through glasses $\checkmark$ High
[77] Smart cane Heavy Low comfort Heavyweight Bulky tablet hangs on neck, heavy cane
[78] Computer-vision-enhanced white cane Not a user-based evaluation: only a technical test performed.
[79] Smart robot No information available.
[80] Smartphone $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ High
TABLE XXV: Practical challenges and operational efficiency - Part II.
Ref. Assistive tool User-friendliness Cost-efficient Weight Comfort for extended use Adjustable fit Fatigue mitigation Portability
Clear instructions Easy to learn
[82] Head-worn camera Easily understood commands 10-minute training Commercially available hardware Heavy (5.5lbs $\approx$ 2.49kg) High
[83] Sensors attached to a white cane Not a user-based evaluation. $\checkmark$ Not a user-based evaluation: only a technical test performed.
[84] No prototype implemented; the proposed approach for real-time global localization was tested with an agent.
[85] Chest-mounted camera Not a user-based evaluation $\checkmark$ Light Not explicitly mentioned. High
[87] Suitcase-shaped robot Improved with intuitive terminology Positive usability scores - Heavy (40lbs $\approx$ 18.14 kg) Physical demand - Heavyweight Bulky and heavy
[88] Robotic cane
[89] Smart cane The focus of the paper is primarily on the technical aspects of the multi-sensory blind guidance system.
[90] Smart glasses Light
[91] Hand-worn device Not a user-based evaluation: only a technical test performed. 52.5 grams
[92] Smart E-glasses Areas for improvement
[93] Google glasses Information not provided; Only a technical test conducted with a laptop PC as a navigator.
[95] A rolling suitcase-shaped device $\checkmark$ Short training session (10-20 minutes) Heavy Weight discomfort Heavy and bulky Space limitations
[96] System testing was not feasible as the system was in the experimental stage.
[97] Smart glasses Not a user-based evaluation: only a technical test performed.
[98] Smartphone $\checkmark$ $\checkmark$ $\checkmark$ Light $\checkmark$ $\checkmark$ $\checkmark$ High
[99] White-cane mounted camera Information not provided; Only a technical test performed.
[100] TurtleBot2 robot (to be replaced by portable device) + wearable sensors Heavy Heavy and bulky Bulky
[101] Paper develops a Reinforcement Learning environment to create a navigation assistant tailored for the BVI community, without user-based evaluation.
[102] Smartphone $\checkmark$ Light $\checkmark$ $\checkmark$ $\checkmark$ High
[103] No prototype implemented; the proposed approach was tested within a research building.
[104] Intelligent autonomous scooter High cost Heavy Bulky
[105] Wearable camera Light High
[106] Computer-vision-enhanced white cane Not a user-based evaluation: only a technical test performed.
[107] Smart cane + Google Tango 225 grams High: light and compact, usable in multi-floor
[108] Helmet-mounted camera + android-based smartphone Not a user-based evaluation: only a technical test performed.
[109] Wearable camera Not a user-based evaluation: only a technical test performed.
[110] iPhone 12 Pro Max Effective language instructions Intuitive navigation process Light - High
[111] Wearable camera Not a user-based evaluation: only a technical test performed.
[112] Person carrier robot (wheelchair) Heavy Bulky
@@ III-C2 Practical challenges and considerations
This section discusses how the approaches proposed by the studies included in our SLR have the potential to improve navigation for BVI people. These studies focused on diverse attributes, such as accurate pose estimation, semantic mapping, sensor fusion, and algorithmic innovations, to improve the quality of BVI navigation. Table XXVI presents the categorization of attributes that contribute to enhancing the mobility and navigation of visually impaired individuals.
To understand the impact of these solutions further, we examined their effectiveness in real-world scenarios. Various localization and mapping techniques have been assessed on the basis of their accuracy, robustness, consideration of dynamic objects, and running time. This evaluation provides insight into the performance of these techniques in practical environments.
In addition, we considered user-based evaluations to gauge user satisfaction and the practical applicability of the proposed system. These evaluations include feedback from actual users, which is crucial for understanding real-world usability and acceptance of assistive technologies.
Furthermore, we provide a detailed overview of the components and technologies used in assistive navigation systems. This helps to understand the practical implementations and innovations proposed by the studies. By examining the system prototypes, we gained insights into the design and functionality of assistive solutions beyond the localization and mapping components. This offers a comprehensive view of how these technologies enhance the mobility and navigation of the visually impaired.
III-D1 Attributes enhancing mobility and navigation
@@
TABLE XXVI: Attributes of SLAM-based navigation systems that contribute to enhancing BVI navigation, along with referenced papers emphasizing each feature.
@@
Precise Localization Using SLAM algorithms, these solutions accurately estimate the position and orientation of visually impaired users. Precise localization is essential for visually impaired navigation systems in order to ensure accurate real-time guidance, obstacle avoidance, and spatial awareness, ultimately enhancing independent and safe mobility. [102, 97, 95, 93, 90, 82, 78, 77, 76, 75, 73, 70, 69, 64, 89, 67]
@@
Real-time assistance This feature ensures that BVI users receive immediate feedback about their environment. Therefore, they can be guided to navigate safely and efficiently, thereby enhancing their overall mobility and independence. [85, 57, 92, 82, 77, 69, 66, 67, 88]
@@
Semantic mapping Semantic mapping involves creating detailed environmental representations that go beyond geometric data and enables visually impaired users to navigate with a deeper understanding of their surroundings. [100, 97, 66, 72, 84]
@@
Both indoor and outdoor navigation Systems that serve both indoor and outdoor settings allow users to transition seamlessly between different environments while receiving consistent support. [56, 65, 104, 85, 108, 67, 111, 68, 74]
@@
Innovative algorithms Innovative algorithms lead to advancements in navigation techniques. These approaches contribute to a more efficient and effective navigation, ultimately improving the overall experience of the visually impaired. [110, 102, 77, 66, 106, 78, 91, 73]
@@
Robotic navigation These solutions employ robotic systems, such as robots, smart canes, scooter, and suitcases, which are equipped with sensors to assist BVI in navigating their environment. [106, 104, 99, 112, 107, 78, 70, 64, 89, 67, 59, 68, 87]